# Region Guided Attention Network for Retinal Vessel Segmentation

### v1

@@ 3.2 Decoder Block
A decoder block in the proposed network consists of a deconvolution operation that upsamples the input feature maps, followed by two depth-wise separable convolution operations similar to those in the encoder blocks. Each decoder block receives input from either the previous decoder block or the bottleneck layer, in addition to features from a corresponding encoder block via a skip connection. The deconvolution operation on the decoder up-samples the feature maps, aligning their spatial dimensions with those of the encoder features. This upsampled feature map is then concatenated with the feature map from the encoder block, preserving fine-grained spatial information critical for precise segmentation. After concatenation, the combined features undergo processing through a series of convolutional layers followed by ReLU activation. These layers refine the upsampled features, enhancing the network’s ability to accurately segment the image. The output of this process is then passed on to the next decoder block, continuing the upsampling and refinement process until the original image resolution is restored. By incorporating skip connections and refining the features through convolutional layers, the decoder blocks effectively reconstruct the segmented image, maintaining spatial accuracy and detail. We have also used a cascaded partial decoder to align the high- and low-level features which are then passed to the main decoders (see Figure 1 partial decoder block). This addition positively impacts the segmentation result and as shown in Table 2, represented by PD (Partial Decoder).
@@ 3.4 Proposed Network Architecture
To explain how the proposed network processes input to obtain the desired segmentation map, some model components are first discussed. Let the model input be defined as $X_{in}$, where $X\in\mathbb{R}^{H\times W\times C}$, and $C_{d}^{n\times n}(*)$ be defined as a separable convolution operation in depth, where $n\times n$ is the kernel size, $\beta_{n}(*)$ be batch normalisation and $P_{max}$ be a non-overlapping max pooling operation of size $2\times 2$. Let $\Re$ be ReLU activation and $Conv$ be the convolutional block given in equation 1. Then the first encoder block processes the input $X_{in}$ and returns two values $S_{1}$ and $p_{1}$ as in equations Eq.(2-3).
@@
Here $S_{1}\in R^{H\times W\times 8}$ and $p_{1}\in R^{\frac{H}{2}\times\frac{W}{2}\times 8}$, where $H$, $W$ and $8$ are the height, width and channel of the feature maps respectively. Now we feed $p_{1}$ to the second encoder block and get $S_{2}\in R^{\frac{H}{2}\times\frac{W}{2}\times 16}$ and $p_{2}\in R^{\frac{H}{4}\times\frac{W}{4}\times 16}$ by repeating equations Eq.(1-3). We repeat the same sequence of steps for the third encoder block passing $p_{2}$ as the input of the layer and get $S_{3}\in R^{\frac{H}{4}\times\frac{W}{4}\times 24}$ and $p_{3}\in R^{\frac{H}{8}\times\frac{W}{8}\times 24}$ as the outputs. By now we are done with three encoder blocks and now we feed $p_{3}$ to the bottleneck layer where Eq.1 is performed and we get the feature map $b\in R^{\frac{H}{8}\times\frac{W}{8}\times 32}$ which will be fed to the decoder blocks. At this point, the feature map obtained from the encoder blocks, $b$, is upsampled in the first decoder block by a deconvolution operation $D^{n\times n}(*)$ where $n$ is the kernel size. After the upsampling operation, the resultant feature map is concatenated with $S_{3}$ and we pass it through the convolution block Eq.1. The resultant features of the first encoder block,$dec_{1}$, is calculated in Eq.4.
@@
The second decoder block operates on $dec_{1}$ and $S_{2}$ as inputs and produces $dec_{2}$ by repeating Eq4 and likewise the third decoder block takes $dec_{2}$ and $S_{1}$ as input and produce $dec_{3}$ using Eq.(4). We then use a partial decoder [33] to aggregate high-level features that have smaller spatial resolution compared to low-level features. For this purpose, we pass $dec_{2}$ through the convolution block (Eq.1) and get $dec_{par}$, as shown in Eq.5.
@@
We run $Pred_{final}$ through a $1\times 1$ convolutional layer followed by sigmoid activation function to obtain the final predicted segmentation mask $Mask$ as given in Eq.8.
@@ 4 Results and Discussion
In this section, we first describe the dataset, followed by implementation details, ablation results and comparative analysis.
@@
CHASE 999$\times$960 28 Train: 20, Test: 8
@@
Table 2: Ablation study on different loss functions with basic UNet, UNet plus region guided IAA Block and with integration of cascaded partial decoder. We get best results using UNet with IAA block, cascaded partial decoder and fewer filters in the encoder and decoder blocks. *with filters (8,16,24,32)
Method Loss Functions Jaccard $\textbf{F}_{\mathbf{1}}$ Recall Precision Acc
UNET IoU 0.6602 0.7951 0.7876 0.8081 0.9648
UNET Dice 0.6621 0.7965 0.8042 0.7936 0.9643
UNET Dice + BCE 0.6584 0.7938 0.7850 0.8080 0.9646
IAA + UNET IoU 0.6569 0.7926 0.7766 0.8160 0.9648
IAA + UNET Dice 0.6725 0.8039 0.8055 0.8094 0.9659
IAA + UNET Dice + BCE 0.6794 0.8087 0.8046 0.8205 0.9671
IAA + PD + UNET Dice + BCE 0.6805 0.8096 0.8106 0.8144 0.9666
IAA + PD + UNET * Dice + BCE 0.6903 0.8166 0.8285 0.8098 0.9677
@@ 4.1 Dataset
We used DRIVE, CHASE_DB1 and STARE datasets for the experiments. DRIVE contains a total of 40 images, 20 of which are used for training and 20 for testing. However, the dataset size is not suitable for deep learning purposes, therefore we augmented the dataset. The augmentation included resizing of all the images from $584\times 565$ to $512\times 512$ and application of horizontal and vertical flips along with $360$ degrees of rotation in the training images and saving the image after every degree of rotation. As a result of the augmentation, the training set size increased to $7260$ from $20$ images. Note that the test images were only resized and no other pre-processing or post-processing was performed on them. Likewise, we applied the same augmentations on the CHASE_DB1 dataset where the 20 training images were resized and augmented while the 8 test images were resized. For the STARE dataset, we used the first $10$ images for training on which we applied horizontal and vertical flips, $360$ degrees rotation and again horizontal and vertical flips on the rotated images. The next $10$ images were used for testing and only resized. The details of the datasets are given in Table 1.
@@ 4.2 Implementation Details
Figure 2: Qualitative results of the proposed method on some sample images from DRIVE dataset. The columns from left to right show the query image, segmentation mask (ground truth), the mask predicted by the model and analytic mask respectively. The green pixels in the analytic mask represent the correctly segmented pixels while the red pixels are the false negatives and the blue pixels are the false positives.
Figure 3: Qualitative results of the proposed method on some sample images from CHASE dataset. The columns from left to right show the query image, segmentation mask (ground truth), the mask predicted by the model and analytic mask respectively. The green pixels in the analytic mask represent the correctly segmented pixels while the red pixels are the false negatives and the blue pixels are the false positives.
Table 3: Comparison of the proposed method with other existing works on the DRIVE dataset. Best results are in bold and dashes indicate unknown results.
Method Recall Precision Acc $\textbf{F}_{\mathbf{1}}$ Params (M)
@@
Proposed Method 0.8285 0.8098 0.9677 0.8166 0.04
Table 4: Performance comparison between the proposed method and some alternative methods on CHASE dataset.
@@
Table 5: Performance comparison of the proposed method with a number of alternatives on STARE dataset.
@@ 4.3 Ablation Study
We conducted a comprehensive series of experiments to explore various configurations, including different loss functions and filter counts within the encoder and decoder blocks. In particular, the investigations revealed that the use of fewer filters than the standard UNET configuration yielded optimal results, as outlined in Table 2. Our meticulous analyses highlighted the superiority of the Dice loss function for retinal vessel segmentation. This loss function effectively addresses class imbalance, leading to improved segmentation accuracy. Furthermore, the integration of region-guided attention mechanism significantly improved network segmentation performance. Specifically, as shown in Table 2, the incorporation of the IAA block increased segmentation performance by approximately $2.7\%$. To effectively integrate global and local contexts, we used a cascaded partial decoder [33]. The global context captures the overall structure of the object, while the local context adds fine-grained details. This dual-context integration ensures a comprehensive and detailed representation of the object. The integration of PD resulted in better connected vessels and hence improved segmentation results. Finally, the reduction in the number of filters is particularly advantageous, as it mitigates the risk of overfitting by limiting model capacity. This approach is especially beneficial for medical image datasets, which are often small in size and pose inherent challenges.
@@ 4.4 Comparison with Existing Works
We conducted a comprehensive evaluation of our proposed model using the DRIVE, STARE, and CHASE datasets, comparing its segmentation performance with existing methods. In Table 3 a detailed comparison is presented that shows the efficacy of the proposed model against the latest approaches on the DRIVE dataset. It is worth noting that the majority of existing work shows a remarkable gap between precision and recall, which means that they do not tackle the class imbalance in terms of background and foreground pixels. The higher precision of the models is due to the excess of background pixels in the image. However, our proposed model effectively overcomes the class imbalance issue, as is evident from the consistent precision and recall results. In Tables4 and Table 5, detailed comparisons of the proposed model with the state-of-the-art approaches on the CHASE and STARE datasets respectively are presented. The results of the experiments demonstrate the notable strength and superiority of the proposed model, both in terms of segmentation performance and computational efficiency. Despite its compact size, with only $0.04$ million learnable parameters, our proposed model outperforms existing models in terms of segmentation accuracy. This underscores the efficiency and applicability of our model on devices with limited memory and computational resources, while still achieving superior segmentation results compared to larger models. The proposed method achieves results superior to those of the state-of-the-art on CHASE and STARE datasets. The region-guided attention block pushes our model to distinguish vessels from the background, and the low number of network parameters helps it evade the overfitting curse commonly encountered when working with medical image datasets. Hence, our model competitively beats the state-of-the-art techniques on DRIVE, CHASE and STARE datasets. Furthermore, Figure 2 provides a qualitative analysis of our model performance on sample query images from the DRIVE dataset. Also, Figure 2 illustrates that our model accurately captures both thick and thin retinal vessels, further confirming its effectiveness in accurately segmenting the structures of the retinal vessels.
@@ 5 Conclusion and Future Work
In this paper, we introduced a lightweight segmentation network with a significantly lower number of parameters (0.04 million), which is composed of the encoder-decoder mechanism along with partial encoder and inverse addition attention blocks for region-guided segmentation tailored specifically for retinal vessels, complemented by an in-depth ablation study focussing on hyperparameter optimisation. The region-guided attention network block focusses on the foreground pushing, whereas cascaded partial decoder aligns the high- and low-level features, and hence improves the performance of the model by $2.7\%$. This comprehensive study provides researchers with valuable information, offering a solid foundation to enhance retinal vessel segmentation without the need for extensive hyperparameter optimisation, thus streamlining future research efforts in this domain.

### v3

@@ 3.2 Decoder Block
A decoder block in the proposed network consists of a deconvolution operation that upsamples the input feature maps, followed by two depth-wise separable convolution operations similar to those in the encoder blocks. Each decoder block receives input from either the previous decoder block or the bottleneck layer, in addition to features from a corresponding encoder block via a skip connection. The deconvolution operation on the decoder up-samples the feature maps, aligning their spatial dimensions with those of the encoder features. This upsampled feature map is then concatenated with the feature map from the encoder block, preserving fine-grained spatial information critical for precise segmentation. After concatenation, the combined features undergo processing through a series of convolutional layers followed by ReLU activation. These layers refine the upsampled features, enhancing the network’s ability to accurately segment the image. The output of this process is then passed on to the next decoder block, continuing the upsampling and refinement process until the original image resolution is restored. By incorporating skip connections and refining the features through convolutional layers, the decoder blocks effectively reconstruct the segmented image, maintaining spatial accuracy and detail. We also used a cascaded partial decoder to align the high- and low-level features that are then passed to the main decoders (see Figure 1 partial decoder block). This addition positively impacts the segmentation result and, as shown in Table 2, represented by PD (Partial Decoder).
@@ 3.4 Proposed Network Architecture
To explain how the proposed network processes the input to obtain the desired segmentation map, some components of the model are first discussed. Let the model input be defined as $X_{in}$, where $X\in\mathbb{R}^{H\times W\times C}$, and $C_{d}^{n\times n}(*)$ be defined as a separable convolution operation in depth, where $n\times n$ is the kernel size, $\beta_{n}(*)$ is batch normalisation and $P_{max}$ be a non-overlapping max pooling operation of size $2\times 2$. Let $\Re$ be the ReLU activation and $Conv$ be the convolutional block given in Equation 1. Then the first encoder block processes the input $X_{in}$ and returns two values $S_{1}$ and $p_{1}$ as in equations Eq.(2-3).
@@
Here $S_{1}\in R^{H\times W\times 8}$ and $p_{1}\in R^{\frac{H}{2}\times\frac{W}{2}\times 8}$, where $H$, $W$ and $8$ are the height, width, and channel of the feature maps respectively. Now we feed $p_{1}$ to the second encoder block and get $S_{2}\in R^{\frac{H}{2}\times\frac{W}{2}\times 16}$ and $p_{2}\in R^{\frac{H}{4}\times\frac{W}{4}\times 16}$ by repeating equations Eq.(1-3). We repeat the same sequence of steps for the third encoder block passing $p_{2}$ as the input of the layer and get $S_{3}\in R^{\frac{H}{4}\times\frac{W}{4}\times 24}$ and $p_{3}\in R^{\frac{H}{8}\times\frac{W}{8}\times 24}$ as the outputs. By now we are done with three encoder blocks and now we feed $p_{3}$ to the bottleneck layer where Eq.1 is performed and we get the feature map $b\in R^{\frac{H}{8}\times\frac{W}{8}\times 32}$ which will be fed to the decoder blocks. At this point, the feature map obtained from the encoder blocks, $b$, is upsampled in the first decoder block by a deconvolution operation $D^{n\times n}(*)$ where $n$ is the kernel size. After the upsampling operation, the resultant feature map is concatenated with $S_{3}$ and we pass it through the convolution block Eq.1. The resultant features of the first encoder block,$dec_{1}$, are calculated in Eq.4.
@@
The second decoder block operates on $dec_{1}$ and $S_{2}$ as inputs and produces $dec_{2}$ by repeating Eq.4 and likewise, the third decoder block takes $dec_{2}$ and $S_{1}$ as input and produces $dec_{3}$ using Eq.(4). We then use a partial decoder [46] to aggregate high-level features that have smaller spatial resolution compared to low-level features. For this purpose, we pass $dec_{2}$ through the convolution block (Eq.1) and get $dec_{par}$, as shown in Eq.5.
@@
We run $Pred_{final}$ through a $1\times 1$ convolutional layer followed by a sigmoid activation function to obtain the final predicted segmentation mask $Mask$ as given in Eq.8.
@@ 4 Results and Discussion
In this section, we first describe the dataset, followed by implementation details, ablation results, and comparative analysis.
@@
CHASEDB1 999$\times$960 28 Train: 20, Test: 8
@@
Table 2: Ablation study on different loss functions with basic UNet, UNet plus region guided IAA Block and with integration of cascaded partial decoder. We get the best results using UNet with an IAA block, a cascaded partial decoder, and fewer filters in the encoder and decoder blocks. *with filters (8,16,24,32)
Method Loss Functions Jaccard $\textbf{F}_{\mathbf{1}}$ Recall Precision Acc p-value
UNET IoU 0.6602 0.7951 0.7876 0.8081 0.9648 0.493
UNET Dice 0.6621 0.7965 0.8042 0.7936 0.9643 0.490
UNET Dice + BCE 0.6584 0.7938 0.7850 0.8080 0.9646 0.481
IAA + UNET IoU 0.6569 0.7926 0.7766 0.8160 0.9648 0.484
IAA + UNET Dice 0.6725 0.8039 0.8055 0.8094 0.9659 0.521
IAA + UNET Dice + BCE 0.6794 0.8087 0.8046 0.8205 0.9671 0.512
IAA + PD + UNET Dice + BCE 0.6805 0.8096 0.8106 0.8144 0.9666 0.514
IAA + PD + UNET * Dice + BCE 0.6903 0.8166 0.8285 0.8098 0.9677 0.512
@@ 4.1 Dataset
We used DRIVE, CHASE_DB1, and STARE datasets for the experiments. DRIVE contains a total of 40 images, 20 of which are used for training and 20 for testing. However, the dataset size is not suitable for deep learning purposes; therefore, we augmented the dataset. The enhancement included resizing all images from $584\times 565$ to $512\times 512$ and the application of horizontal and vertical flips along with $360$ degrees of rotation in the training images and saving the image after every degree of rotation. As a result of the augmentation, the training set size increased to $7260$ from $20$ images. Note that the test images were only resized and no other pre-processing or post-processing was performed on them. Likewise, we applied the same augmentations on the CHASE_DB1 dataset where the 20 training images were resized and augmented while the 8 test images were resized. For the STARE dataset, we used the first $10$ images for training on which we applied horizontal and vertical flips, $360$ degrees rotation, and again horizontal and vertical flips on the rotated images. The next $10$ images were used for testing and only resized. The details of the datasets are given in Table 1.
@@ 4.2 Implementation Details
Figure 2: Qualitative results of the proposed method on some sample images from the DRIVE [47] dataset. The columns from left to right show the query image, segmentation mask (ground truth), and the mask predicted by the model and analytic mask respectively. The green pixels in the analytic mask represent the correctly segmented pixels while the red pixels are the false negatives and the blue pixels are the false positives.
Figure 3: Qualitative results of the proposed method on some sample images from the CHASEDB1 [48] dataset. The columns from left to right show the query image, segmentation mask (ground truth), and the mask predicted by the model and analytic mask respectively. The green pixels in the analytic mask represent the correctly segmented pixels while the red pixels are the false negatives and the blue pixels are the false positives.
Table 3: Comparison of the proposed method with other existing works on the DRIVE [47] dataset. The best results are in bold, and dashes indicate unknown results. Some works do not compute the $F_{1}$ Score for their network performance, hence there is ’-’ in the table.
Method Sensitivity Specificity Acc $\textbf{F}_{\mathbf{1}}$ Params (M)
@@
Proposed Method 0.8285 0.9822 0.9677 0.8166 0.04
Table 4: Performance comparison between the proposed method and some alternative methods on the CHASEDB1 [48] dataset.
@@
Table 5: Performance comparison of the proposed method with a number of alternatives on the STARE [56] dataset.
@@ 4.3 Ablation Study
We conducted a comprehensive series of experiments to explore various configurations, including different loss functions and filter counts within the encoder and decoder blocks. In particular, the investigations revealed that the use of fewer filters than the standard UNET configuration yielded optimal results, as outlined in Table 2. Our meticulous analyses highlighted the superiority of the Dice loss function for retinal vessel segmentation. This loss function effectively addresses class imbalance, leading to improved segmentation accuracy. Furthermore, the integration of a region-guided attention mechanism significantly improved network segmentation performance. Specifically, as shown in Table 2, the incorporation of the IAA block increased segmentation performance by approximately $2.7\%$. To effectively integrate global and local contexts, we used a cascaded partial decoder [46]. The global context captures the overall structure of the object, while the local context adds fine-grained details. This dual-context integration ensures a comprehensive and detailed representation of the object. The integration of PD resulted in better connected vessels and hence improved segmentation results. Finally, reducing the number of filters is particularly advantageous, as it mitigates the risk of overfitting by limiting the capacity of the model. This approach is especially beneficial for medical image datasets, which are often small in size and pose inherent challenges.
@@ 4.4 Comparison with Existing Works
We conducted a comprehensive evaluation of our proposed model using the DRIVE [47], STARE [56], and CHASEDB1 [48] datasets, comparing its segmentation performance with existing methods. In Table 3 a detailed comparison is presented that shows the efficacy of the proposed model against the latest approaches on the DRIVE dataset. It is worth noting that the majority of existing work shows a remarkable gap between sensitivity and specificity, which is due to the class imbalance in terms of background and foreground pixels. The higher specificity of the models is due to the excess of background pixels in the image. However, our proposed model achieves greater sensitivity and specificity, which shows that our model segments the foreground and the background pixels better than the existing work. In addition, our model achieves the highest accuracy among existing works, showing that we capture the most retinal vessels accurately. In Tables4 and Table 5, detailed comparisons of the proposed model with the state-of-the-art approaches on the CHASEDB1 and STARE datasets are presented, respectively. The results of the experiments demonstrate the notable strength and superiority of the proposed model, both in terms of segmentation performance and computational efficiency. Despite its compact size, with only $0.04$ million learnable parameters, our proposed model outperforms existing models in terms of segmentation accuracy. This underscores the efficiency and applicability of our model on devices with limited memory and computational resources while still achieving superior segmentation results compared to larger models. The proposed method achieves results superior to those of the state-of-the-art CHASEDB1 and STARE datasets. The region-guided attention block pushes our model to distinguish vessels from the background, and the low number of network parameters helps it avoid the overfitting curse commonly encountered when working with medical image datasets. Hence, our model competitively beats the state-of-the-art techniques on DRIVE, CHASE, and STARE datasets.
4.5 Qualitative Results
We present some of the qualitative results obtained by our proposed method. Figure 2 provides a qualitative analysis of our model performance on sample query images from the DRIVE dataset. Furthermore, Figure 2 illustrates that our model accurately captures both thick and thin retinal vessels, further confirming its effectiveness in accurately segmenting the structures of the retinal vessels. We showcase the qualitative performance of the proposed method on the CHASEDB1 dataset [48] in Figure 3. We distinguish the correctly labeled pixels, false positives, and false negatives in the last column of the figure. The qualitative results, too, approve the efficiency of the proposed model in segmenting retinal vessels.
@@ 5 Conclusion and Future Work
In this paper, we introduce a lightweight segmentation network with a significantly lower number of parameters (0.04 million), which is composed of the encoder-decoder mechanism along with partial decoder and inverse addition attention blocks for region-guided segmentation tailored specifically for retinal vessels, complemented by an in-depth ablation study focused on hyperparameter optimisation. The region-guided attention block focuses on the foreground push, whereas the cascaded partial decoder aligns the high and low level features, and hence improves the performance of the model by $2.7\%$. This comprehensive study provides researchers with valuable information, offering a solid foundation to enhance retinal vessel segmentation without the need for extensive hyperparameter optimisation. Thus, streamlining future research efforts in this domain.